2014 daily meeting minutes

In doubt about an n_TOF acronym? Use the n_TOF Glossary page!

6 August, Wednesday

Target: Everything is fine.

Beam line 1: no news

Detectors in EAR1: - SiMon: acquired data during last evening and last night. Seen gamma-flash and signals;

- SI2D: are working, there is just some anomaly in one digitizer that seems to give a wrong input to the SiMon. 6 Strips are not connected to give channels to mMegas. - mMegas: high noise probably related to the preamplifier. U5 converter was installed. - C6D6: calibrations data were taken this morning with an 88Y gamma source.

NB: RP went there to check and the experimental area was not clean, we also miss two tools. Pleas be careful to not bring any tools outside the area (it's a class-A lab) and clean up your thing when you finish to operate in the area.

Beam line 2: - new lead spheres were installed, still missing the first ones.

- Vacuum: the big leak was now fixed and the levels are around 3-5 10^-2. There is a problem in the upper part: the segment between the upper valve and the detectors has to be replaced. The upper valve still cannot be open by remote control, sensors and valves are under investigation.

Detectors in EAR2: - C6D6: measurements of the background and gamma-flash done.

- Simon2: will be installed this afternoon to start taking signals tonight.

Plan: Run with all the strips of SI2D in order to have the profile of the beam (once the PTB is installed the detector cannot be used in this position -> maybe we'll move it after Simon).

-- MauroLicata - 6 Aug 2014

5 August, Tuesday

Beam Line 2: - Vacuum: lower part and beam dump ok (at the level of the pump). Problem in the part in the experimental area (1.2 10^-1 mbar), under investigation. The valve cannot still be opened, but we can have beam with the valve closed.

- Lead spheres will be put in place this afternoon.

Laser alignment system: ok in EAR1, has still to be fixed in EAR2

Beam Line 1: - Magnet: problem solved, the piquet came and reboot the power supply. NB since for now wham we enter in the experimental area (= break the search) the magnet is turned off automatically, we need to put the power supply in stand by mode BEFORE asking for the access mode!

- The first door to enter in EAR1 has to stay open when we are inside.

DAQ: there is an online monitor to see signals online. Still problems in the new routines to generate root and dat files.

Detectors: - mMegas: tomorrow morning the 235U will be put inside the detector

- PTB: once put in operation with Nolte we can keep it till about the end of August.

We will have beam from this afternoon till tomorrow morning.

On Call phone number: 71018

-- FedericaMingrone - 5 Aug 2014

4 August, Monday

DAQ: During the week end it has been noticed a problem with the trigger in one digitizer. It had been disconnected with the consequent loss of 2 channels.

- PKUP signals is not properly attenuated, so data recorded during the weekend are useless. Problem to be solved this afternoon.

- It is not possible to create properly nor the dst neither the root files, so it is impossible to check if detectors are working fine.

Need time to sort out the problem (also need to define the priorities Alvaro should work on), so there will be no beam tonight.

Beam line 1: - Magnet: it switches off when the current increases above 290 A (it should work at 600 A). Need to operate on the power supply, but a special procedure is needed (it is not possible to operate on the magnet when Access mode is ON). It should be necessary perhaps one day.

EAR1: - Detectors: PTB is arrived today, its mechanical support is work in progress and should operate next week.

- Iron shielding will be ready for the end next week -> need to coordinate with PTB people to try to extend the use of the detector for some time in order to operate with the final configuration of EAR1.

Beam Line 2: - Vacuum: the leak has been removed, the problem was due to an incorrect joint.

- The polyethylene is in, as well as the lead balls.

- The beam dump will be completed this afternoon.

The beam-line should be ready for tomorrow!!

NB: Thursday 7 is fixed a ventilation test. No access both in EAR1 and EAR2 for all the day.

The new start of n_TOF has appeared in the CERN Bulletin, please find below the link where you can find the article.

“THE FIRST NEUTRON BEAM HITS EAR2” http://cds.cern.ch/journal/CERNBulletin/2014/32/News%20Articles/?ln=en

-- MauroLicata - 4 Aug 2014

1 August, Friday

DAQ: Yesterday crashed after a few hours. The problem is only in the module 2, stream 2, is not working, and 2 channels are off.

EAR1: Everything is fine.

- Available 26 channels for Simon 2D, 4 channels for MGAS and 4 for Simon.

- Simon 2D: Data are been acquired with a gamma source. This source will be removed before 6 pm and the beam will be requested. The source put for C6D6 has been already removed.

Beam line 2: lower part aligned. Collimators and polyethylene will probably be set up before 5 pm.

The beam is expected after 6 pm. The intensity of the beam will be the same as the last week (about 1/10)

-- MauroLicata - 1 Aug 2014

31 July, Thursday

Beam Line 1: - Filter station: cabled, temporary manual control (8 switcher, move right -> out, move left -> in). The rack is close to the cooling station.

- Access issues. Magnet: changing the interlock problem -> easy to solve technically but problems with the bureaucracy. Search box in EAR1: the second key will be removed, but for the moment we will need to call the CCC to do the search in EAR1. The "beam on" warning will be activated on Christmas.

EAR1: Simon2d has all the DAQ channels required. Sandbox.MicroMegas is again installed (4 channel in DAQ).

Beam Line 2: - Vacuum: need to dismount everything, they will work on it next week.

We will start taking data in EAR1 at night and during the weekends

Alberto Mengoni prepared a glossary with all the terminology of n_TOF and how to use abbreviations and conventions: https://twiki.cern.ch/twiki/bin/view/NTOF/UnifiedGlossary. Take advantage of it!

NB: if you enjoy a coffee in the control room, please pay for it!

-- FedericaMingrone - 31 Jul 2014

30 July, Wednesday

Target & Cooling Station: Everything is fine. The conductivity is the same as the 2012 and a new corrosion monitor is available. Ready to receive beam.

EAR1: - Beam Line: tomorrow there will be a meeting with Access team to communicate our requests. Filter box has been cabled this morning (still waiting to install control system). NB Vacuum people will need 1 day to access all the beam line and test it.

- Experimental area: an 88Y source is put in place to see signals with C6D6 detectors (only background was not sufficient to commission the DAQ).

EAR2: - Vacuum: NO leakage in the lower part (below magnet). Upper part tested with a new turbo-molecular pump but the limit still remain at 50 mbar -> need to investigate the line to find the leak.

- Experimental area: SiMon removed to install the polyethylene blocks. Ventilation problem (too high Δp, opening the door was very hard) solved for the moment, ventilation people will need anyway 1 day without anyone in the area and with everything in place. On 5th august there will be interventions for grounding.

Conclusion: - there will be no beam for sure until tomorrow morning. Anyway, the PS will be ready to give us full intensity by the end of the week.

- an article on the new start of the facility will appear on the CERN Courier on friday (we provided them the picture of the first signal seen in EAR2).

-- MauroLicata - 30 Jul 2014

29 July, Tuesday

Target & Cooling Station: ok

EAR1: - Beam line: ask to the PS access team to solve the switching off of the magnet every time we enter the area, this afternoon they'll have a meeting to decide what to do. NB: also to do the search is required to call CCC to put the area in "search mode". The filter box is going to be cabled tomorrow but it hasn't any remote control: it should be implemented to be operative on monday 11 August. If it is not possible, we can have a temporary local solution.

- Experimental area: next tuesday (5th August) the laser system will be installed in both experimental area => NO BEAM. Detectors in EAR1: SiMon -> good, SiMon2D -> good, need all the channels, C6D6 -> 1 Bicron + 1 L6D6 on the floor to commission the DAQ, Sandbox.MicroMegas -> dismounted to compare as a reference with Sandbox.MicroMegas in EAR2, will be put back tomorrow. The 235U foil will be mounted on tuesday 5/8.

Sources: need to find a strategy for calibration sources, i.e. alpha unsealed or Cf, which are expensive to buy but difficult to borrow: a dedicated meeting will take place in the afternoon.

EAR2: - Beam line: now we have a leak detector to investigate the beam line. Lower part (below the magnet): a metallic joint not properly mounted was found, at 5 pm we will know if the problem is due to this piece. Upper part (above the magnet): could have small leakages, need to understand if they are fixable without removing all the collimators. Tomorrow they are going to install the second part of the polyethylene block. The beam dump will be completed mid august. !!! NB !!! ANY INTERVENTION IN THE AREA HAS TO BE COMMUNICATED TO ALL THE RESPONSIBLE PERSONS OF THE DETECTORS MOUNTED.

- Experimental area: tomorrow they will come to install the mobile phone network (the landline is already working). The cables for the Radiation Monitor (RadMon) are already placed, tomorrow before 4 pm they will test the connections. The intervention to put cable trays and grounding is foreseen for one of our beam off days, i.e. 5/8, 8/8, 11/8. Detectors in EAR2: SiMon2 -> yesterday the remaining 3 Si detectors were mounted together with the Mylar+LiF foil (now the detector is in the final configuration, so pay attention when playing around). Need to think a final and "professional" solution for the light tight of the kapton window, for the moment a temporary Al foil is put on top. Sandbox.MicroMegas: the detector was cleaned in the lab and repaired (there was a short circuit from dirt and mesh), it will be out back tomorrow. Diamond: they were dismounted and could be put back next week when needed. C6D6: 1 Bicron + 1 L6D6 mounted to commission the DAQ.

DAQ is now working!! there are two potential items which are under further investigations. Tomorrow we will have a test with C6D6+source signals.

-- FedericaMingrone - 29 Jul 2014

28 July, Monday

Target & Cooling Station:

At the target station we had a problem during weekend. The pressure in the cooling circuit was increasing due to the existence of air in this borated/moderator circuit. This must have affected the concentration of boron and temperature in the circuit . Because of this, the beam was stopped on Sunday at 4pm and the problem has been solved this morning (Monday) by purging the circuit. The values of pressure, air and conductivity should be checked during the shifts. A "shift checklist" must be created in order to monitoring all these values. Alvaro is going to include some alarms in the DAQ if these values cross the thresholds.

EAR1: - Beam line: The magnet went off during the weekend. Apparently, it is now interlocked with the access to any EAR. Simone will try to get this to the previous situation so that we don;t have to re-start the magnet every time we enter the EAR. Filter box is in place, but not operational yet.

- Experimental area: Alignment of not done (the detector are fixed on the flanges). The laser will be installed next week (Tuesday).

EAR2: -Beam line: There was a leak in the bottom of the vacuum tube. The vacuum tube wasn't correctly alignment. Now they are pumping vacuum's tube section by section in order to find and solve the problem. The valve in the EAR-2 must be properly connected and interlocked. The last part of the collimator is being machine and is expected to be installed on Wednesday; while the remaining stainless steel balls are coming in the first week of August.

- Experimental area: There new HV cables ready for install and label in EAR2.

* Detectors:* - SiMon-1 (EAR 1): It's working fine. Nothing to report. - MGAS (EAR 1): Seems working, but data need to be checked. - SILI-2D (EAR 1): It's working fine. Nothing to report. They plug-in an MCA and obtained a pulse height spectrum. - SiMon-2 (EAR 2): It's working fine. Nothing to report. - MGAS (EAR 2): It seems that there is a shortcut between different pads/anode/mesh. - Diamonds (EAR 2): It's working fine.Nothing to report.

DAQ: It's not working yet, but Alavaro believes it will be ready on Wednesday.

Beam expected for next Wednesday afternoon.

Javier Balibrea

25 July, Friday

Dear All,

This is to inform you that today, 25 July 2014 at 17:30, after the 18 months shutdown (LS1), the first 20 GeV proton beam pulse has been sent to the n_TOF target. The n_TOF target is now feeding neutrons to our experimental area (EAR-1) at 185m as well as to the new vertical experimental area, EAR-2, at 20m from the spallation source.

It is a great moment for the n_TOF Collaboration to which we arrived only with the full support and help from the CERN management, the EN, GS, TE and BE CERN departments, the n_TOF team at CERN, the CERN Radiation Protection Service, and of course from all the members and institutions of the n_TOF Collaboration.

We can tell you already that we have seen signals from neutron induced reactions and from the arrival of the γ-flash in both EAR-1 and EAR-2 with several of our existing and new detectors, the next few days and weeks will be certainly very exciting at the n_TOF Control room. The commissioning of the n_TOF EAR-2 and of the rest of the facility will now proceed as planned and you will be updated with the progress through the daily meeting minutes and other reports.

Attached a picture of the toast after receiving the first beam!

Happy and excited for the future ahead of us,

Enrico Chiaveri
Spokesperson of the n_TOF Collaboration

CERN-EN/STI CH-1211 Geneva 23 Switzerland tel. +41 22 7676189 gsm. +41 76 4870091 fax +41 22 7668213

24 July, Thursday

At the target station, operations proceeded as planned for the restart of the cooling and moderation systems. The relevant parameters of both circuits, can be accessed online from the DAQ WebSite (see in particular the Interactive Plot feature at the bottom of the page by clicking on "JSPlot"). The ventilation system monitors revealed some air leaking which prevents the pressure difference of the target station controlled area with the outside to reach the threshold value.The places in which the leaks occur have been identified and appropriate measures are in place to solve the problem.

The filter box has been mounted and the whole beam line leading to EAR-1 is already in vacuum.

In EAR-1, the SiMON chamber with the Silicon detectors have been mounted and the beam line is now closed. The Micromegas detector will be mounted tomorrow morning (though it is not yet sure it can be operated immediately). The SiMON-2D detectors have been mounted and the signals tested.

In EAR-1 the Δp is now -65 Pa, allowing to consider it effectively a class-A work sector.

The PE and B4C material components for the core part of the second collimator of the vertical beam line leading to EAR-2 arrived this morning at CERN. However, some adjustment needs to be made before mounting, therefore it was decided to install only 1/2 of the PE section. The line has been then completed with the beam shutter, the SiMON-2 chamber and the upper part from the beam dump down to about 1.5m from the lower side window (some photos of the lower part of the beam line are posted here). The second Micromegas detector will be mounted tomorrow morning for a first test. The diamond detector will be mounted only in EAR-2.

The vacuum condition in the whole vertical line is being tested.

It is foreseen to have everything ready for the PS picket inspection by 10h00, tomorrow. Beam for n_TOF should be delivered by 12:00 noon.

-- AlbertoMengoni - 24 Jul 2014

23 July, Wednesday

At the target station, the borated water is being loaded into the moderator circuit. A new system for avoiding boron crystallization has been installed and will be tested during the run. The conditions for accepting beam (by the target station) will be reached tomorrow evening.

The filter box of the EAR-1 beam line mounting is being completed. 5 positions for filters are retained (see below the corrected table shown in yesterday's meeting minutes). The pumping for vacuum in the whole beam line leading to EAR-1 will be started as soon as the filter box mounting is finished.

The ventilation system in EAR-1 will be tested tomorrow. If de-pressurization conditions will not be immediately met (as required for a class-A work sector), the run can start with EAR-1 as a standard class-C work sector with no impact on the startup operations.

The SiMON Li-foil will be mounted today and the beam line inside EAR-1 will be closed as soon as possible. The micromegas detector will be mounted independently (not connected directly to the SiMON chamber). The SiMON-2D detector cabling and full testing is going on and will be completed by tomorrow.

The shielding of the beam dump, on top of the EAR-2 building, is in progress. Both the iron and concrete blocks will be in place within today.The beam line with the detector components will be mounted tomorrow. SiMON-2 will start up with one Silicon detector for initial testing. Micromegas detecor for EAR-2 will be mounted but will start operating later on, after the gas supply and cabling is completed.

Diamond detectors will be tested, under the new experimental conditions, both in EAR-1 as well as in EAR-2.

IMPORTANT: the cleaning of EAR-2 will take place tomorrow (THU) from 15h00 to 18h00. No access to the area will be allowed during this time period.

Meanwhile, a problem of access and beam-stop control conflicts surfaced, due to the different access conditions for EAR-1 and EAR-2. A discussion with the department DSO is ongoing and a clarification should be reached by tomorrow.

A DAQ system full test will be performed today.

In conclusion, all things considered, the present conditions should allow us to ask the PS control center to prepare to deliver the proton beam to n_TOF by tomorrow evening. Beam for actual startup measurements should start effectively on Friday.

-- AlbertoMengoni - 23 Jul 2014

22 July, Tuesday

Beam for n_TOF is still planned to start on 24 July (THU).

The filter box is being mounted in the beam line leading to EAR-1. 6 positions 5 positions for filters are retained. They are (see here):

slot material thickness [mm] note
1 Al 50 -
2 Pb 20 -
3 Pb 10 -
4 Cd 0.5 with 1 mm Al support frame
5 W 0.8 sandwich
Ag 0.5
Mo 1.0
Co 0.25

In EAR-1, the SiMON chamber has been mounted as well as the rest of the beam line in the area. Micromegas mounting will follow. A new silicon monitor/detector system will be in the beam, mostly for testing. This is a 2D silicon strip detector with a LiF layer in front, mounted on a C-fiber backing. The detector will be mounted in front of the latest part of the beam line in EAR-1 and will not interfere with the SiMON + Micromegas detectors.

* Special thanks due to Daniela Macina, Carlos Guerrero and Jorge Lerendegui who took care of the cabling in EAR-1, assuring the availability of all the needed daq channels for the run startup.

As planned, the permanent magnet has been moved in and the beam line is now completed up to the floor level in EAR-2. The first pipe and valve-shutter will be mounted tomorrow. The beam line mounting will then proceed with the SiMON-2 chamber and continue with the rest, from the top down.

-- AlbertoMengoni - 22 Jul 2014

21 July, Monday

Beam for n_TOF is planned to start on 24 July (THU).

The mounting of the various components of the beam-line to EAR-1 is in progress. From the target area up to the last window before the wall of EAR-1 (tunnel side), the beam line will be completed by tomorrow afternoon. The filter-box will be mounted tomorrow, with controls for the filters to be implemented later on. Vacuum operations and final checks will be performed on Wednesday. Alignment of the EAR-1 components will also be performed on Wednesday.

Detectors in EAR-1 for the first beam will be SiMON and Micromegas (loaded with B and Li. U will be in later on).

Mounting of the various components of the beam-line for EAR-2 is proceeding. The (permanent) magnet is ready to be moved on the line. The core of the second collimator (B4C + polyethylene) will be mounted within two weeks from start as part of the material needed is not yet at CERN. The shielding around the second collimator will be completed after the full amount of material (iron spheres) will be delivered (1/5 of the total already set in place). The beam-dump on top of EAR-2 is being assembled. Weather should improve to allow the mounting of the outer iron and concrete shielding layers by tomorrow or, at the latest, on Wednesday. Also for the beam-dump, additional B4C/PE and lead components will be installed as soon as the material will arrive at CERN.

A similar set of detectors as in EAR-1 will be in place in EAR-2, i.e. SiMON-2 and a new set of Micromegas.

The start-up configuration will allow to begin the commissioning of the two n_TOF beam lines. In particular, alignment and beam profile will be performed in EAR-1, after the modifications of the beam line in the second collimator area. For the vertical beam-line leading to EAR-2, the first beam will allow to measure and monitor the radiation field in several points, mostly for radiation protection purposes. Commissioning will then proceed with the measurement of the beam profile, flux and background conditions for the various detection setups.

-- AlbertoMengoni - 21 Jul 2014

-- AlbertoMengoni - 07 Aug 2014

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2020-08-20 - TWikiAdminUser
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox/SandboxArchive All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback